Grade 12Physical Chemistry

0.1 Molal aqueous solution of an electrolyte AB3 is 90% ionized. The boiling point of solution at 1atm is Kb(H2O) =0.52 K Kg mol^-1

Ranjith
8 Years agoGrade 12

1 Answer

Abhiraj Pathak
8 Years ago
∆Tb = Tb -Tbo = i.Kb.m
AB3 A3+ + 3B- 
n = 4
 
a = i-1/n-1
0.9 = i-1/4-1
i = 3.7
 
Tbo = 373K
 
Thus,
Tb = (3.7*0.52*0.1)+373 = 373.192K